Skip to end of metadata
Go to start of metadata

Slow progress today.  Split into same 3 crews, except two crews also had 2 each INEGI climbers.  Each INEGI crew installed 2 towers:

  • tnw09: up, with solar power.  Connects to network, but connection not yet built to get out of the valley.
  • tnw11: up, installed, but all 3 RMYs report continuous bad data.  This is very odd, since they worked in the ops center and (some) other RMYs work elsewhere in the experiment, so we don't think we're doing something systematically wrong.
  • rne06: up, installed, on network, but it appears that the 20m sonic anemometer head was not plugged in – an easy fix.
  • rne07: installed (including soils), but DC power connection is not live (even when a power supply was swapped with tnw11).  Not clear what is going on.

The 3rd crew prepped v06 and v07 at the base of the tower and measured remaining guy wire tensions.  We also installed soil sensors at tse02.

More prep work also was done at the ops center.

The question is where to go from here...



  1. Additional info on tnw11: After Ted and Gary installed it with Ricardo and Joao Paolo.  The 20m sonic is actually at 20.54m, because we did not successfully explain that the center of the sonic needed to be at 20m.  The 10m was installed much closer to 10m.  The 2m is actually a little higher because they moved it up to avoid clamping over the lacing like had been done for the December installs.

    The bulgin power cable is currently strung through the trees above the road, and it's a little bit of a stretch to get it to the tower.  Probably it needs a more permanent solution, such as stringing cord or cable from the power pole to the tower and hanging the bulgin from that, or burying conduit maybe?


  2. Additional info on tnw09: Ted and Gary visited at the end of the day to add the 5m bulgin cable to the soil mote.  Ted tried to connect to the Ubiquiti but couldn't.  Since it worked for Steve, something seems to have changed.